A man is at a distance of 6m from a bus. The bus begins to move with a constant acceleration of 3m/s2. In order to catch the bus, the minimum speed with which the man should run towards the bus is :
(a) 2 m/s
(b) 4m/s
© 6m/s
(d) 8m/s

Grade:11

1 Answers

Khimraj
5 years ago
let speed is v
then to catch bus
6 + ½*3*t2 = v*t
3t2 -2vt + 12 = 0

Now the logic is simple, the passenger will catch the bus after a certain time. So the passenger can catch the bus only when time is a real number. So the quadratic equation must give t as a real number.

For that the discriminant (D) of the equation must be greater than or equal to zero.

4v2 – 4*3*12 >=0

for min velocity

v2 = 36

v = 6 m/s

Hope it clears. If you Like answer then please approve the  answer.